### v1

# V2Xum-LLM: Cross-Modal Video Summarization with Temporal Prompt Instruction Tuning

Abstract
Video summarization aims to create short, accurate, and cohesive summaries of longer videos. Despite the existence of various video summarization datasets, a notable limitation is their limited amount of source videos, which hampers the effective fine-tuning of advanced large vision-language models (VLMs). Additionally, most existing datasets are created for video-to-video summarization, overlooking the contemporary need for multimodal video content summarization. Recent efforts have been made to expand from unimodal to multimodal video summarization, categorizing the task into three sub-tasks based on the summary’s modality: video-to-video (V2V), video-to-text (V2T), and a combination of video and text summarization (V2VT). However, the textual summaries in previous multimodal datasets are inadequate. To address these issues, we introduce Instruct-V2Xum, a cross-modal video summarization dataset featuring 30,000 diverse videos sourced from YouTube, with lengths ranging from 40 to 940 seconds and an average summarization ratio of 16.39%. Each video summary in Instruct-V2Xum is paired with a textual summary that references specific frame indexes, facilitating the generation of aligned video and textual summaries. In addition, we propose a new video summarization framework named V2Xum-LLM. V2Xum-LLM, specifically V2Xum-LLaMA in this study, is the first framework that unifies different video summarization tasks into one large language model’s (LLM) text decoder and achieves task-controllable video summarization with temporal prompts and task instructions. Experiments show that V2Xum-LLaMA outperforms strong baseline models on multiple video summarization tasks. Furthermore, we propose an enhanced evaluation metric for V2V and V2VT summarization tasks.
11footnotetext: Equal Contribution.22footnotetext: Corresponding author.
1 Introduction
Figure 1: Illustration of cross-modal video summarization.
The interest in sharing life experiences has surged in recent years, making video the most informative and diverse visual medium on social media platforms. This trend has led to significant demands for a variety of video and language understanding tasks, such as video captioning [79], video question answering [82, 78], moment retrieval [32], and video summarization [14, 61]. Video summarization (V2V) provides an efficient way for humans to obtain key information from a long video. This process entails selecting the most significant information from a video and condensing it into a shorter form, all while maintaining the essence of the original content. Beyond extensive research in V2V summarization, there are a few recent explorations of V2T and V2VT summarization. Most notably, VideoXum [37] seeks to broaden the modality of video summaries to include text summaries. It utilizes the dense captions from ActivityNetCap [31] videos as text summaries and annotates the corresponding video segments as video summaries. Other datasets for video summarization include TVSum [61], SumMe [14], QFVS [57], MED Summaries [53], and so on. While it is expected that powerful LLMs can help improve video summarization, the insufficient number of source videos may not be able to support the robust fine-tuning of LLMs to perform this task since finetuning a large model with a limited number of training examples is prone to overfitting [20, 21]. In addition, the data compiled in VideoXum cannot be considered genuine summaries because the dense captions of ActivityNetCap contain a significant amount of redundant information. More recently, Shot2Story20K [15] collects 20k video-text data that enables the robust finetuning of LLMs, but it only supports video-to-text summarization. To address these issues, we propose Instruct-V2Xum, a new large-scale cross-model video summarization dataset that contains 30k open domain videos, partitioned as 25,000 in the training set, 1,000 in the validation set, and 4,000 in the test set. In Instruct-V2Xum, we obtain the source videos from YouTube using the video list provided by InternVid [74]. The methodology of video summarization parallels that of extractive text summarization, where the objective is to isolate the pivotal frames or sentences from the source videos or documents, respectively. Extractive text summarization is a foundational task in the NLP field, with numerous LLMs setting the benchmark in this domain [85, 26, 17]. Inspired by this, we first extract frames from the source videos and employ LLaVA-1.5-7B [39] to generate detailed captions for each frame. Then we take all the frame captions as a document to perform extractive document summarization using GPT-4 [1]. This approach enables us to obtain both video summaries and their corresponding textual summaries. Finally, the extracted text summaries are further refined into a grammar-fluent short paragraph by GPT-4.
Numerous visual instruction tuning-based methods have been proposed for video-language understanding [22, 70, 36]. These models can process the long videos for general video-language understanding and reasoning tasks such as video question answering, video captioning, and so on. Recently, there have been some attempts for fine-grained video moments understanding [22] or video-language temporal grounding [38] using large VLMs. However, these approaches, which typically process video frames as sequential images for a frozen visual encoder and train an LLM decoder for identifying video moment boundaries, are not well-suited for dense temporal prediction in video summarization tasks, particularly in the V2VT tasks. Furthermore, most existing models require large-scale data to train new parameters added to pre-trained VLMs [38, 36, 44]. This requirement significantly limits their practicality in scenarios where only a limited number of training examples are available. To address these problems, we design a new temporal prompt instruction tuning framework – V2Xum-LLaMA. In V2Xum-LLaMA, we unify different modalities of video summary generation into one LLM decoder. This framework stands out by removing the dependence on task-specific layers that were required for video summarization in earlier VLM-based approaches. The main advantages of this method are that it enables the effective adaptation of the learned knowledge and the powerful capabilities of the pretrained language models to dense video temporal and content understanding. All the pretrained parameters of the VLMs are reused, and the model takes interleaved video frames and natural language temporal prompts as input to facilitate end-to-end model training. As video temporal prediction is performed by using language models, there is a challenge for calculating the correlation for V2V evaluation, since the language model-predicted video summaries are the discrete frame indexes. To overcome this challenge, we propose a solution in section 5 to calculate the scores for language model-predicted video summaries. Furthermore, we also provide an analysis of the existing video summarization tasks and propose $F_{CLIP}$ and $Cross-F_{CLIP}$, the enhanced evaluation metrics for V2V and V2VT summarization tasks.
In summary, our main contributions are as follows:
  • We propose a novel and flexible cross-modal video summarization framework – V2Xum-LLaMA, where we unify different video summarization tasks into one pre-trained language decoder without introducing the randomly initialized task-specific heads. The proposed framework takes interleaved video frames and natural language temporal prompts as input, which facilitates the pre-trained VLMs to process long sequences in an end-to-end manner. Our proposed method outperforms all the selected strong baseline models on existing mainstream V2V, V2T, and V2VT benchmarks.
  • To address the challenge of insufficient video-language data for effectively fine-tuning large VLMs for various video summarization tasks, we build a new instruction following dataset named Instruct-V2Xum for cross-model video summarization. This dataset comprises 30k diverse videos sourced from YouTube, with durations spanning from 40 to 940 seconds. The proposed dataset enables VLMs to generate modality-controllable video summaries with the task prompt. The experiments validate the rationality of our proposed dataset.
  • We provide a comprehensive analysis of various existing video summarization tasks from the perspectives of data, method, and evaluation. Building on this analysis, we propose $F_{CLIP}$ and $Cross-F_{CLIP}$, an enhanced evaluation metric for V2V and V2VT summarization tasks. Experimental results indicate that our proposed evaluation metrics are highly consistent with the traditional evaluation metrics including F1, Spearman correlation, and Kendall correlation.
2 Related Work
2.1 Video Summarization
Video summarization is the process of distilling salient information from video content by analyzing its structure and temporal redundancies. Traditional video summarization, also known as video-to-video summarization, typically generates a condensed version of the original video, comprising selected frames [41, 13, 71, 8], shots [25, 9, 88], or segments [66, 30]. These models are commonly trained using supervised learning approaches, with reinforcement learning methods like policy gradient [77], optimizing for diversity and representativeness in the summarized output [91], gaining popularity. The standard datasets for video summarization tasks include SumMe [14] and TVSum [61], which are widely used for benchmarking purposes. In recent years, cross-modal video summarization [11, 16, 24] has emerged as an area of interest, incorporating additional modalities such as audio, speech, subtitles, and captions. This approach leverages multimodal models to create more comprehensive summaries. Video-to-text summarization [50, 4] is an evolving field that aims to generate descriptive paragraphs in natural language that encapsulate video content. VideoXum [37] makes progress in this direction by repurposing the ActivityNetCap [31] dataset and employing the vision-language model BLIP-2 [33] for both video-to-video and video-to-text summarization. However, it is important to note that the data collected in [37] may not constitute genuine summaries, as the dense captions from the ActivityNetCap dataset often contain significant redundant information.
2.2 Video Temporal Understanding
Beyond video summarization, there has been an increasing interest in temporal understanding tasks for video analysis. Temporal localization plays a pivotal role in comprehending long-form or untrimmed videos by associating specific segments with their corresponding semantics. Research has primarily concentrated on tasks such as temporal video grounding [42, 75], dense video captioning [72, 81], video highlight detection [32, 28], step localization [58], and procedure planning [2]. Despite significant progress, models designed for action temporal localization and highlight detection often rely on predefined labels [83], which hampers a comprehensive understanding. To address these limitations, research has focused on generic event boundary detection [59, 73, 67] and dense video captioning [31], utilizing datasets that provide event descriptions [37]. Nevertheless, Most existing models still rely on regression for temporal predictions, necessitating separate heads for captioning and regression tasks [72, 86, 66]. Recent advancements in Large Language Models (LLMs) have introduced a paradigm shift by enabling the direct use of natural language to specify temporal locations in videos, thus offering a more intuitive approach to video understanding.
2.3 Large Language Models
In recent years, Large Language Models (LLMs) have witnessed rapid advancements [1, 68, 69, 89]. With pretraining on extensive corpora from the Internet, LLMs acquire substantial knowledge, enabling powerful zero-shot and in-context learning capabilities [1, 76, 29, 19]. Efforts have been increasingly directed toward leveraging LLMs for multimodal tasks [43, 60, 63]. Techniques such as vision-language alignment and adapter fine-tuning are employed to integrate LLMs into the multimodal domain. These methods align the visual features extracted by visual encoders with the input token space of LLMs [39]. Typically, the parameters of LLMs are frozen to retain their existing capabilities although LoRA [18] fine-tuning is occasionally applied to adapt to data with varying distributions or to acquire new behaviors. Utilizing this approach, several studies [70, 44, 36, 84] have successfully employed LLMs for video understanding tasks, referred to as Vid-LLMs [64]. However, current research primarily concentrates on general video understanding tasks, such as video question-answering (QA) and video captioning, with less emphasis on temporal information. Recent works [22, 55, 65] have started exploring LLMs’ potential in temporal grounding and localization, highlighting the untapped potential of LLMs in video temporal understanding.
Table 1: Comparison with existing video-to-video summarization and video-to-text summarization datasets. “V2V”, “V2T”, and “V2VT” denote whether the datasets support video-to-video summarization, video-to-text summarization, or both tasks, respectively. “Instruction” denotes whether the dataset supports video-text instruction tuning. M and S stand for manual and model synthesized, respectively.
Dataset Domain # Videos Annotation V2V-Sum V2T-Sum V2VT-Sum Instruction
MSVD [3] Open 1,970 M
YouCook [5] Cooking 88 M
MSR-VTT [79] Open 7,180 M
UCF101 [62] Open 13,320 M
ActivityNetCap [31] Activities 20,000 M
Shot2Story20k [15] Open 20,000 M+S
SumMe [14] Events, holidays, sports 25 M
TVSum [61] News, documentaries, vlogs 50 M
OVP [6] Documentaries, lectures 50 M
VSUMM [7] Cartoons, news, commercials, shows 50 M
EDUVSUM [12] Letures 98 M
LoL [10] Matches of League of Legends 218 M
Ads-1K [66] Commercials 1,041 M+S
VideoXum [37] Activities 14,001 M
Instruct-V2Xum Open 30,000 M+S
Figure 2: Distribution of the duration of source videos (in seconds) the summaries length for both video and text in Instruct-V2Xum.
3 The Instruct-V2Xum Dataset
In this section, we introduce the details of data collection and provide the quantity and quality analysis for Instruct-V2Xum, which is an instruction-tuning dataset for cross-modal Video summarization.
3.1 Data Curation
3.1.1 Frame Captioning and Extractive Summarization.
The source videos are sampled from InternVid [74]. We first filter the raw videos according to the duration of the videos and their aesthetic scores. We obtain 30k for video and text summaries generation. Initially, we extract frames at a rate of 1 FPS, followed by the transformation of these sampled video frames into their detailed textual representations using LLaVA-1.5-7B [39]. Upon obtaining the text representations of the video frames, we utilize GPT-4 to perform extractive document summarization. Finally, the frame captions from the text summaries are converted into both video and text summaries.
3.1.2 Text Summarization Refinement.
To further reduce the redundancy of the text summaries, we utilize BERT score [90] to filter out the frame text representations that are similar to other summary frames. Here, we set the threshold to 0.93. The filtered video frame captions’ indexes serve as the video summaries for the source videos. Then, we employ GPT-4 to further compress and rewrite the video summaries to be shorter and more grammar-fluent.
3.2 Quantity Analysis
Here we provide the statistical results for our collected data. Figure 2 illustrates the distribution of the duration of source videos and the summaries length for both video and text in Instruct-V2Xum. We also provide more statistical results in the appendix. The average duration of the source videos is 183 seconds. The average length of the text summaries is 239 tokens. The average length of the video summaries is 30 frames. The average compression ratio is 16.39%.
3.3 Quality Analysis
Since the text summaries are synthesized by LLMs, we analyze the quality of the generated summaries from the perspective of grammar fluency and commonsense plausibility. For the grammar-fluency evaluation, we leverage a grammar-check model [46] that assigns high scores to correct texts grammatically. For the commonsense plausibility evaluation, we utilize Vera [40], a plausibility estimation model, to identify the nonsensical bias. The results of these evaluations are presented in the appendix. Overall, our data demonstrates reasonable scores in both grammar and Vera evaluations.
Figure 3: The architecture of our proposed V2Xum-LLaMA.
4 Methodology
4.1 Overview
This section presents our proposed unified cross-modal video summarization framework V2Xum-LLaMA. Our method is distinguished by its utilization of temporal prompt encoding and temporal-aware decoding to condense and summarize video content effectively. As depicted in Fig 3, our model architecture comprises three key components: video frames are encoded through a vision encoder, subsequently combined with temporal prompts, and finally processed by a decoder-only LLM. Utilizing LLaMA as the decoder, our system is capable of generating both video-to-video and video-to-text summarizations. Temporal Prompt Encoding associates video frames with corresponding timestamps, embedding positional information into each frame. These encodings are then merged with specific summarization instructions, and the summaries are produced through temporal-aware decoding via LLaMA. Our methodology allows for modality control over the generated summaries, based on the provided instructions, producing either video or text summarizations, or a combination thereof. This innovative encoding-decoding mechanism enhances the model’s understanding of the video timeline and improves the summarization output’s relevance and accuracy.
4.2 Interleaved Video and Temporal Prompt Encoding
The temporal prompt mechanism is designed to bind the visual tokens with the corresponding timestamps at the frame level to inject positional information for each frame. To this end, we first encode each video frame $f_{i}$ using pretrained CLIP [54] encoder $E_{v}$. The frames’ encoding $v_{i}=E_{v}(f_{i})$ are involved in the visual tokens sequence $V=\{v_{1},v_{2},...,v_{L}\}$, where $L$ is the number of sampled frames.
We then bind temporal prompts with each visual token. The temporal prompts are tokenized zero-padded numbers in natural language format like “00”, “06”, “12”, “99”, etc., indicated by $T=\{t_{1},t_{2},...,t_{L}\}$. They are inserted into the visual token sequence $V$ to form a new sequence with interleaved visual tokens and temporal prompts:
$S=\{t_{1},v_{1},t_{2},v_{2},...,t_{L},v_{L}\}.$
Compared to the original visual token sequence, the temporal prompted sequence can better capture the relations between the timestamps and the visual semantics.
4.3 Temporal-Aware Decoding
We use decoder-only LLM, LLaMA [69] as the decoder to generate video-to-video summarization $A^{v}$, video-to-text summarization $A^{t}$, or both $A^{b}$. The type of generated summarization is controlled by the instructions $I^{t}$, $I^{v}$, or $I^{b}$ that request the model to generate video summarization, text summarization, or both, respectively. With the instruction $I$ and temporal prompted sequence $S$, the output is given by the temporal-aware decoding:
$A^{x}=LLM(S,I^{x}),~{}x\in\{v,t,b\}.$
The video-to-video summarization $A^{v}$ is defined as a sequence of temporal tokens sequence, i.e., $A^{t}=\{t_{i}\}^{M}_{i=1}$, which is a subset of temporal prompts sequence $T$ and $M\leq L$. The video-to-text summarization is a summarized caption in a natural language format. In our implementation shown as Fig. 3, the video-to-text summarization also contains frame referring temporal tokens, i.e.:
$A^{t}=\{...,w_{i-1},t_{j},w_{i+1},w_{i+2}...\},$
where the temporal token $t_{j}$ is bind with the words or sentences that consist of the summarization of visual content and can be further extracted from $A^{t}$ to get the video-to-video summarization $A^{v}$. This is called temporal-aware decoding. An example is shown in Fig. 3. When decoding, the temporal tokens can be decoded together with the text summaries and represent the temporal position where the described visual content occurred in the input video.
4.4 Task-Controllable Video Summarization
As mentioned before, the types of summarization can be controlled by the instructions $I^{x}$, and we will further introduce how to utilize the instruction following the capability of LLM to realize the modality controllable video summarization.
We use instructions like “Please generate a/BOTH VIDEO/TEXT summarization for this video.” as the instructions and tokenize them with the same text tokenizer that tokenizes the temporal prompts.
During the instruction tuning, we freeze the parameters of the vision encoder and update the vision adapter, an MLP that aligns the visual semantic with the input token space of the LLM, and the language decoder to force it to generate temporal indexes.
5 Experiments
Table 2: Comparison Results on the VideoXum dataset. “TSH-Free” means the model is task-specific-head-free; “B-4” denotes the BLEU-4; “M” denotes the METEOR metric; “R-L” refers to the ROUGE-L metric; “C” represents the CIDEr metric for consensus-based evaluation. For all evaluation metrics, higher scores indicate better performance.
Method Cross- Modal LLM- Based TSH- Free V2T V2V V2VT
B-4 M R-L C F1 Spearman Kendall FCLIP Cross-FCLIP
DENSE [31] 1.6 8.9 - - - - - - -
DVC-D-A [34] 1.7 9.3 - - - - - - -
Bi-LSTM+TempoAttn [93] 2.1 10.0 - - - - - - -
Masked Transformer [93] 2.8 11.1 - - - - - - -
Support-Set [52] 1.5 6.9 17.8 3.2 - - - - -
Frozen-BLIP [33] 0.0 0.4 1.4 0.0 16.1 0.011 0.008 - -
Vid2Seq-HCY [81] 2.3 8.2 19.0 7.6 24.2 - - 0.888 0.214
Vid2Seq-HC [81] 2.7 8.5 19.8 8.4 24.5 - - 0.892 0.217
Vid2Seq-HCV [81] 2.7 8.4 19.8 8.3 25.1 - - 0.899 0.200
VSUM-BLIP [37] - - - - 21.7 0.207 0.131 - -
TSUM-BLIP [37] 5.6 11.8 24.9 20.9 - - - - -
VTSUM-BLIP [37] 5.8 12.2 25.1 23.1 23.5 0.258 0.196 0.894 0.247
V2Xum-LLaMA-7B (ours) 5.8 12.3 26.3 26.9 29.0 0.298 0.204 0.931 0.253
V2Xum-LLaMA-13B (ours) 5.7 12.3 26.2 25.3 31.6 0.276 0.200 0.957 0.251
Human - - 5.2 14.7 25.7 24.2 33.8 0.305 0.336 0.944 0.256
Table 3: Comparison results on the TVSum and SumMe datasets.
Method TVSum SumMe
Spearman Kendall Spearman Kendall
dppLSTM [87] 0.055 0.042 - -
DSN [91] 0.020 0.026 - -
Sumgraph [51] 0.138 0.094 - -
CLIP-it [48] 0.147 0.108 0.120 0.109
TL;DW [47] 0.167 0.143 0.128 0.111
iPTNet [27] 0.174 0.148 0.131 0.114
A2Summ [17] 0.178 0.150 0.143 0.121
Standard ranker [56] 0.230 0.176 0.014 0.011
VSUM-BLIP [37] 0.261 0.200 0.365 0.268
V2Xum-LLaMA 0.293 0.222 0.378 0.296
5.1 Baseline Models
Our V2Xum-LLaMA model, including both 7B and 13B versions, is benchmarked against various models adept at video-to-video (V2V) summarization, video-to-text (V2T) summarization, or both tasks on the VideoXum dataset. Baseline models include LLM-based approaches such as Frozen-BLIP [33], VSUM-BLIP [37], TSUM-BLIP [37], and VTSUM-BLIP [37]. We compare with task-specific-head-free (TSH-Free) models like DENSE [31], DVC-D-A [34], Bi-LSTM+TempoAttn [93], Masked Transformer [93], and Support-Set [52], which do not rely on regression-based timestamp prediction with extra task-specific heads. Additionally, on the classical TVSum [61] and SumMe [14] datasets, we compare our 7B version V2Xum-LLaMA with the following V2V summarization methods: dppLSTM [87], DSN [91], Sumgraph [51], CLIP-it [48], TL;DW [47], iPTNet [27], A2Summ [17], Standard ranker [56], and VSUM-BLIP [37]. We also evaluated Vid2Seq [81] models, initially pre-trained on the HowTo100M [45], VidChapter-7M [80], YouCook2 [92], and ViTT [23] datasets, and subsequently fine-tuned on the VideoXum dataset to establish baseline models for our experiments.
5.2 Evaluation Metrics
We introduce new CLIP-based evaluation metrics for V2V and V2VT summarization evaluation. While the F1 score is a common metric for V2V summarization tasks, the process of video summarization annotation is highly subjective, leading to considerable variance among human annotators [49]. The traditional F1 score, which compares predicted video frames directly with the ground truth, fails to account for semantically similar frames that are close in time but not exactly matching, thus potentially undervaluing accurate summaries. To mitigate this, we introduce the $F_{CLIP}$ metric for V2V summarization evaluations, which is designed to recognize and reward semantic similarities between predicted and ground truth frames even when they are not identical. And we also propose the $Cross-F_{CLIP}$ metric for the V2VT summarization tasks. Unlike the VT-CLIPScore metric used in VideoXum—which calculates the average cross-modal CLIP score as an indicator of semantic alignment between predicted video and text summaries—our $Cross-F_{CLIP}$ calculates the average $F_{CLIP}$ scores between the predicted video summaries and the ground truth text summaries, as well as between the predicted text summaries and the ground truth video summaries. This approach aims to provide a more nuanced evaluation of summarization tasks by acknowledging the importance of semantic content alignment across modalities. For a reference video summary $v$ and predicted video summary $\hat{v}$, the recall, precision, and F1 scores are:
$R_{CLIP}(v,\hat{v})=\frac{1}{|v|}\sum_{v_{i}\in v}\max_{\hat{v}_{j}\in\hat{v}{}}\mathbf{v}_{i}^{\top}\mathbf{\hat{v}}_{j}$ (1)
$P_{CLIP}(v,\hat{v})=\frac{1}{|\hat{v}|}\sum_{\hat{v}_{j}\in\hat{v}}\max_{v_{i}\in v}\mathbf{v}_{i}^{\top}\mathbf{\hat{v}}_{j}$ (2)
$F_{CLIP}(v,\hat{v})=2\frac{P_{CLIP}\cdot R_{CLIP}}{P_{CLIP}+R_{CLIP}}$ (3)
For the $Cross-F_{CLIP}$, given a reference video and text summary $v$ and $t$ and the predicted video summary $\hat{v}$ and text summary $\hat{t}$:
$Cross-F_{CLIP}(v,\hat{v},t,\hat{t})=\frac{F_{CLIP}(v,\hat{t})+F_{CLIP}(\hat{v},t)}{2}$ (4)
Given that the cosine similarity values range from -1 to 1, we adjust the similarity scores by applying the operation $\max(\cos(\mathbf{v},\mathbf{\hat{v}}),0)$. This ensures that only non-negative similarity scores are considered in our analysis.
5.3 Implementation Details
We use CLIP ViT-L/14@336 as the vision encoder and Vicuna-v1.5-7B/13B as the large language model. The vision adapter is pre-trained by vision-text alignment with image-text pairs from LCS-558K [39] dataset. After being processed by the Vicuna-v1.5 tokenizer, each temporal prompt takes four tokens. To facilitate model training, we normalize the video length into 100 via downsampling. For each setting, we train the V2Xum-LLaMA for 5 epochs with a learning rate of 1e-4 on 8$\times$×NVIDIA A100 GPUs.
5.4 Quantitative Results
To verify the effectiveness of our proposed method, we conduct experiments on multiple video summarization benchmarks, including cross-modal video summarization and video summarization tasks. For cross-modal video summarization, we evaluate our method using the VideoXum dataset [37] along with our newly proposed V2Xum dataset. In the realm of V2V summarization, we utilized TVSum [61] and SumMe [14] as benchmarks for our evaluation.
5.4.1 Cross-Modal Video Summarization
The experimental results of VideoXum are shown in Table 2. It can be summarized from the table that our proposed V2Xum-LLaMA model outperforms all the baseline models. Specifically, in the task of V2V summarization, V2Xum-LLaMA achieves a 8.1% higher F1-Score than VTSUM-BLIP, alongside significant improvements in both Spearman and Kendall correlation metrics. In addition, V2Xum-LLaMA-13B achieves higher evaluation scores on the V2V summarization task than V2Xum-LLaMA-7B. On the contrary, V2Xum-LLaMA-7B performs better on V2T summarization. We attribute this result to the increased complexity of the V2T summarization compared to V2V summarization. Additionally, the VideoXum training set comprises only 8,000 examples, a quantity insufficient for effectively training models with large language decoders.
We present the performance of various models on our newly proposed Instruct-V2Xum dataset, with the results detailed in Table 4. The results indicate that the models are well-adapted to the dataset, exhibiting sound performance. Moreover, V2Xum-LLaMA-13B outshines V2Xum-LLaMA-7B in V2T summarization, which we believe is due to the larger language models benefiting from the increased volume of training data.
Table 4: Comparison results on the Instruct-V2Xum test set.
Method V2T V2V V2TV
B-4 M R-L C F1 FCLIP Cross-FCLIP
Vid2Seq-HC [81] 3.8 6.1 22.6 0.4 23.0 80.5 16.1
Vid2Seq-HCY [81] 3.7 6.2 22.4 0.5 24.7 81.3 16.0
Vid2Seq-HCV [81] 3.6 6.2 22.5 0.4 25.1 81.5 16.3
V2Xum-LLaMA-7B 6.8 15.8 26.9 0.9 31.7 95.5 23.1
V2Xum-LLaMA-13B 6.7 15.8 27.0 0.8 31.3 95.3 23.0
5.4.2 Video-to-Video Summarization
We evaluate the capability of V2Xum-LLaMA on VideoXum, TVSum, SumMe, and Instruct-V2Xum datasets. The results are shown in Table 2 and Table 3. It can be concluded from the empirical results that the unified video summarization using the language decoders in V2Xum-LLaMA can effectively perform traditional V2V summarization tasks. V2Xum-LLaMA outperforms all the previous methods that relied on task-specific regression heads for generating video summaries. This result indicates that LLMs with temporal prompts are capable of performing fine-grained video temporal understanding. The results presented in Table 4 affirm the validity of our proposed V2Xum dataset. It demonstrates that the model can properly fit the data.
5.5 Ablation Study
To further evaluate the efficacy of our proposed V2Xum-LLaMA framework and to underscore the importance of augmenting the training dataset, we conduct an ablation study on V2V and V2VT summarization tasks. First, we explore the performance enhancements attributed to the introduction of the temporal prompt mechanism, with findings detailed in Table 5. The data reveal that integrating the temporal prompt into the V2Xum-LLaMA framework yields superior results in both V2V and V2T summarization tasks compared to models lacking this feature, which indicates the temporal prompt benefits language models for fine-grained dense video temporal understanding. Second, to analyze the impact of concurrently generating video and text summaries on the performance of V2V and V2T summarization tasks, we conduct ablation studies. The results, as illustrated in Table 5, suggest that the simultaneous generation of video and text summaries enhances the performance in the V2V summarization task. Third, we also compare the performance of models training with more data (Instruct-V2Xum + VideoXum) and the models finetuned without more data. The results indicate that instruction data can effectively improve models’ performance, especially for the V2T summarization task. Furthermore, we conduct a comparison between models equipped with a pre-trained vision adapter and those without to highlight the critical importance of pre-training the vision adapter in enhancing performance. The results indicate that a well-pre-trained visual prompt adapter enhances the models’ performance for V2V and V2T tasks. Finally, we provide a comparison of parameter-efficient finetuning (PEFT) and full parameter fine-tuning. The results are presented in Table 5. The results demonstrate that full parameter fine-tuning outperforms PEFT in performance.
Table 5: Ablation Study of our V2Xum-LLaMA (7B) on the VideoXum dataset.
Method V2T V2V V2VT
BLEU-4 METEOR ROUGE-L CIDEr F1-Score Spearman Kendall FCLIP Cross-FCLIP
V2Xum-LLaMA 5.8 12.3 26.3 26.9 29.0 0.298 0.204 0.931 0.253
w/o simultaneous VT-Sum 5.6 12.2 25.6 25.9 25.1 0.249 0.203 0.926 0.251
w/o Instruct-V2Xum 4.9 12.0 24.3 21.6 23.1 0.260 0.191 0.921 0.252
w/o fully fine-tuning 4.5 11.7 24.7 22.8 23.4 0.222 0.175 0.915 0.250
w/o temporal prompts 4.4 11.7 24.4 21.2 23.9 0.258 0.192 0.910 0.249
w/o pretrained adapter 3.1 11.1 21.9 9.5 3.7 - - - -
6 Discussions on Video Summarization: Data, Method, and Evaluation
In this section, we discuss the existing vision summarization tasks from the perspective of data, method, and evaluation. As mentioned before, current existing video summarization datasets, including V2V and V2VT video summarization datasets, contain few training examples and cannot support training large-scale deep neural networks. The TVSum dataset comprises merely 50 YouTube videos; the SumMe dataset includes only 25 personal videos sourced from YouTube, while the QFVS dataset provides 135 video-query training samples. Additionally, VideoXum, the first cross-modal video summarization dataset, provides only 8k training examples. Our experimental results reveal that it is insufficient for training 13B models for the V2T summarization task. This issue is one of the significant drawbacks of current existing datasets. To address this problem, we collect more videos from YouTube and generate the corresponding cross-modal summaries for the videos using GPT-4 and propose a large-scale cross-modal video summarization dataset.
A conventional approach to V2V summarization involves training a regression head to assign an importance score to each frame, ranking frames based on these scores, and selecting the top Kframes to evaluate performance using metrics like the F1 score or Kendall/Spearman correlation against the ground truth. However, as demand for cross-modal video summarization grows, this method is inadequate for multimodal video summarization. Recent studies have begun to explore the use of language models for generating temporal and spatial references in videos [22, 35, 55], demonstrating the viability of using language models to generate the video interval indexes. In addition, leveraging large language models’ text decoders for both V2V and V2T summarization tasks is a logical step. Therefore, in this study, we investigate how to prompt large VLMs to understand the fine-grained video content along with temporal information, how to maximally leverage the powerful capability of content understanding and reasoning of large language models, and how to achieve task controllability via natural language instructions. To that end, we first propose the temporal prompt mechanism, and then design the visual encoder that takes the interleaved video frames, temporal, and language as input. Instead of using the regression head to perform V2V summarization, we unify different video summarization tasks into one language decoder.
Evaluating video summaries poses a significant challenge, primarily because the criteria for quality are inherently subjective, vary across different viewers and even fluctuate over time. This subjectivity, coupled with the limited availability of evaluation videos and annotations, further exacerbates the ambiguity in assessing video summary quality [49]. The traditional F1 score, designed to directly compare predicted video frames with ground truth, ignores the nuances of semantically similar frames that, while temporally proximate, do not exactly match. This oversight can lead to the undervaluation of otherwise accurate summaries. To address this, we design new CLIP-based F scores for evaluating V2V and V2VT summarization tasks. Unlike the traditional F1 score, which relies on exact matches for its precision and recall calculations, $F_{CLIP}$ evaluates the V2V summarization from the perspective of semantic similarity. This approach allows for a more comprehensive yet meaningful evaluation that recognizes the importance of semantic accuracy over mere frame-by-frame accuracy. To evaluate the alignment of the generated video and text summaries, VideoXum employs an approach that calculates the average CLIP score across video frames and text summaries. This method involves computing vectors for all frames and sentences, applying mean pooling to these vectors to represent the video and text summaries, and then calculating the cosine similarity between them. However, this approach evaluates the video and text summaries as unified entities and thus neglects the detailed alignment at the sentence level with specific video frames. In contrast, our proposed $Cross-F_{CLIP}$ adopts a greedy matching strategy to optimize the similarity score between individual video frames and corresponding sentences, ensuring a more granular and accurate alignment.
7 Conclusion
In this study, we address the deficiencies in current video summarization datasets including the insufficient number of training examples and the insufficient evaluation of video summarization by building a new large-scale cross-model video summarization dataset Instruct-V2Xum and designing the improved video summarization evaluation metrics. We also propose a novel temporal prompt instruction tuning method – V2Xum-LLM that unifies the generation of different modalities of video summaries into the text decoder of VLMs. This method does not rely on task-specific heads to perform video summarization tasks and enables the models to accept the interleaved long video and language input sequences. In addition, our proposed method also supports modality controllable summary generation through language instructions. The experimental results show the effectiveness of our proposed method.
Appendix A Importance Score Prediction with LLM Logits
Since our V2Xum-LLaMA directly utilizes natural language to represent the temporal positions, it is hard to obtain the corresponding importance scores that are usually adopted to calculate some conventional video summarization evaluation metrics, such as Spearman’s $\rho$ and Kendall’s $\tau$. Therefore, we use the logits from LLM that decode the temporal positions (numbers) as the importance scores.
Specifically, we first find the indices of the positions that are numbers in the output of V2Xum-LLaMA. Then, the corresponding logits are processed by the Softmax function to obtain the probabilities of the tens place and ones place:
$p_{i}(tens)$ $=p_{i}(x_{t}|x_{<t})=\text{SoftMax}(\text{logit}_{t}),$
$p_{i}(ones)$ $=p_{i}(x_{t+1}|x_{<t+1})=\text{SoftMax}(\text{logit}_{t+1}),$
where $t$ is the index in the output sequence that the ten places of the temporal tokens occur. Next, we approximate $p(\text{tens},\text{ones})$ with the product of $p(\text{tens})$ and $p(\text{ones})$:
$p_{i}(tens,ones)\approx p_{i}(tens)\times p_{i}(ones).$
Finally, we compute the average of all the $p(tens,ones)$ generated from the prediction in natural language format and treat it as important scores for calculating evaluation metrics:
$imp\_score=\frac{1}{M}\sum_{i=i}^{M}p_{i}(tens,ones).$
Appendix B Quality Analysis
We analyze the quality of GPT-Synthesised data from the perspective of grammar fluency and commonsense plausibility. For the grammar-fluency evaluation, we leverage a grammar-check model [46] that assigns high scores to correct texts grammatically. For the commonsense plausibility evaluation, we utilize Vera [40], a plausibility estimation model, to identify the nonsensical bias. The results are presented in Figure 4 and Figure 5. We can conclude that our dataset with the refinement process achieves a satisfactory score.
Figure 4: Comparison of the Vera scores for the dataset before (left) and after (right) refinement, with higher scores indicating better results.
Figure 5: Comparison of the Grammar scores for the dataset before (left) and after (right) refinement, with higher scores indicating better results.
Appendix C Example Data from Instruct-V2Xum
In this section, we show some data points sampled from the Instruct-V2Xum training set in Figure 6.
Figure 6: Some videos with video summaries and text summaries annotations from our Instruct-V2Xum dataset.